OpenStudy (anonymous):

If cosx=2/3 and x is in Q4, then sin(x/2)=?

myininaya (myininaya):

do you know any identities that involve a half angle?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

sin(x/2)=+/- sqrt((1-cosx)/2)

myininaya (myininaya):

ok replace cos(x) with what is given

myininaya (myininaya):

then follow order of operations yes the same order of operations learned in beginning algebra

OpenStudy (anonymous):

I'm lost in your first step... Replace what exactly with what?

myininaya (myininaya):

cos(x)=2/3 so that tells me i can replace cos(x) with 2/3

myininaya (myininaya):

let's see if he can get the answer

OpenStudy (anonymous):

So...replacing cos(x) with 2/3, I am given sqrt((1-2/3)/2) Is this correct so far?

myininaya (myininaya):

and for prettiness i'm going to rewrite our identity \[\sin(\frac{x}{2})=\pm \sqrt{\frac{1}{2}(1-\cos(x))} \\ \sin(\frac{x}{2})=\pm \sqrt{\frac{1}{2}(1-\frac{2}{3})}\]

myininaya (myininaya):

now can you follow the order of operations to simplify this?

OpenStudy (anonymous):

Yes, it's clear to me now.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (bradely):

+ or -sqrt((1-2/3)/2) =+ or -sqrt((3-2/3)/2) =+ or -1/sqrt (6)